5. A 0.400 mg initial sample of radioactive iodine–123, used to treat thyroid cancer, decreases to 0.113 mg after 24.0 hrs. Determine the half–life of iodine–123. 


1
Expert's answer
2022-04-26T08:20:02-0400

m = m02-t/T

m/m0 = 0.400 mg / 0.113 mg = 1/4

2-24/T = 1/4

2-24/T = 2-2

-24/T = -2

T = 12 hrs